OpenStudy (anonymous):

Write the equation in standard form y=-2/5x+1

OpenStudy (anonymous):

a.2x+5y=5 b.-5x=2y+5 c.x=-2/5y+1 d.-5y=2x+5

OpenStudy (texaschic101):

standard form is Ax + By = C y = -2/5x + 1 --- multiply the entire equation by the LCD, 5, to get rid of fractions 5y = -2x + 5 -- add 2x to both sides 2x + 5y = 5